EXCITED STATE PERTURBATIONS OF 7-AZAINDOLE MEDIATED THROUGH MICRO-SOLVATION.^{a}

\underline{JUSTIN W. YOUNG}, and DAVID W. PRATT, Department of Chemistry, University of Pittsburgh, Pittsburgh, PA 15260.

7-azaindole readily forms water clusters when seeded into a molecular beam. Interestingly, the addition of water considerably perturbs the excited state of 7-azaindole: the electronic origin of 7-azaindole(H_{2}O) is shifted drastically to lower energy, 1290 cm^{−1}, relative to the origin of the bare molecule; furthermore, recent evidence indicates that 7azaindole(H_{2}O) is of L_{a} character, contrary to the bare molecule.^{b} In this work, high resolution electronic spectra of 7-azaindole(H_{2}O) were measured in the presence of an external electric field , as a means to study water induced changes in the electronic distribution of 7-azaindole. Furthermore, the L_{a}/L_{b} character of the excited state of 7-azaindole(H_{2}O) was examined by recording similar spectra of isotopically labeled species.

aWork supported by NSF(CHE-0911117)

bT. B. C. Vu, I. Kalkman, L. W. Meerts, Y. Svartsov, C. Jacoby, and M. Schmitt, J. Chem. Phys. 128, 214311 (2008).